# Krylov Complexity in early universe

### v2

@@ 6.5 The evolution of $r_{k}(\eta)$ and $\phi_{k}(\eta)$ in open system’s method
where prime denotes the varying with respect to the conformal time $\eta$. The same as the closed system, we change variable as $y=\log_{10}a$, the evolution equation of $r_{k}$ and $\phi_{k}$ will become as,
@@
During the calculations, we accounted for the background of $\phi$ using conformal time $\eta$, as demonstrated by the various potentials in Eq. (23). Consequently, $\mu_{2}$ and $\mu_{1}$ also vary with respect to conformal time, since they include contributions from the potential and the scale factor $a$.
@@
Fig. 8 clearly illustrates the evolution of $r_{k}$ across various epochs, including inflation, radiation domination (RD), and matter domination (MD). It is evident that the different potentials produce similar trends for $r_{k}$. Unlike in Fig. 3, there is no oscillation of $r_{k}$ once dissipation effects are considered. Additionally, the trend of $r_{k}$ varies, particularly during the MD phase, where the evolution consistently shows an increase. These differences will result in distinctive evolutions of Krylov complexity and Krylov entropy.
6.6 Krylov complexity and Krylov entropy with the approach of open system
In this section, we will continue use the wave function (73) to explore the Krylov complexity and Krylov entropy. All of these calculation are exactly the same with our previous study Li:2024kfm , the only differences is that we need to use $r_{k}$ to take place of $\eta$ in the formulas for Krylov complexity and Krylov entropy whose formulas can be expressed as following:
@@ 6.6 Krylov complexity and Krylov entropy with the approach of open system
where we have defined $A=1+2\mu_{2}\tanh r_{k}+(\mu_{2}^{2}-|1-\mu_{1}^{2}|)\tanh^{2}r_{k}$ for simplicity. It straightforwardly demonstrates the leading order of Krylov complexity (76) and Krylov entropy (77) will nicely recover into the case of Eq. (40) and Eq. (41) under the weak dissipative approximation, namely, $\mu_{2}\ll 1$.
Fig. 9 illustrates the evolution of Krylov complexity during inflation, radiation domination (RD), and matter domination (MD). During inflation, we observe that Krylov complexity increases exponentially, which is consistent with the closed system’s method. However, the trends in RD and MD differ from those shown in Fig. 4, where the complexity tends to stabilize around certain constant values, oscillating around these values in both RD and MD. In the case of an open system, it is evident that Krylov complexity decreases during RD. For MD, Krylov complexity initially peaks as the scale factor $a$ increases and then declines to stabilize at constant values. Additionally, it is noteworthy that the order of Krylov complexity in an open system is significantly lower than in a closed system, even in RD and MD. Our numerical analysis shows that the term $\rm sech(r_{k})$ in Eq. (76) leads to this outcome, regardless of whether the system is weakly dissipative or not. We have demonstrated that both RD and MD can be approximated as weakly dissipative systems under the condition $\mu<1$. Recalling the meaning of Krylov complexity, it describes the operator growth in a many-body system and our early universe is a perfect chaotic system due to $b_{n}=\alpha n$. Our numeric reveals that the operator growth is much smaller compared with the case of a closed system, which is consistent with reality since the dissipation will lead to faster decreasing of operator growth compared with that without the dissipation effects.
Once one understands the role of dissipation effects, it becomes evident that the order of Krylov entropy in an open system is lower compared to that in a closed system, as illustrated in Figs 10 and 5. The evolution trend is particularly distinctive for RD and MD, as shown in Figure 10, where the Krylov entropy initially peaks before stabilizing at constant values. In the case of MD, the Krylov entropy initially holds steady at certain values before decreasing to some constant level. Overall, the background conditions and dissipative effects have a significant impact on both Krylov complexity and Krylov entropy.
@@
In this section, we utilized the wave function presented in Eq. (73) to investigate Krylov complexity and Krylov entropy. By comparing our findings to those from a closed system, we observed that the dissipative effects not only influence the overall evolution of the system but also reduce the order of Krylov complexity and Krylov entropy. Since Krylov complexity reflects the growth of operators—a fundamental quantum property of a dynamical system—we can conclude that, when the growth of the operator is diminished in the open system compared to the closed system, the dissipative effects lead to a more rapid decoherence-like behavior. Meanwhile, we also could see that the various potentials will almost produce the same trend of Krylov complexity and Krylov entropy in different epochs.
@@ 7 Summary and outlook
The definition of Krylov complexity is clear and unambiguous, especially when compared to circuit complexity. Its uniqueness eliminates the need for introducing a parametric manifold to illustrate the concept of complexity. Additionally, Krylov complexity characterizes the growth of operations within a dynamical system, and the order of Krylov complexity can indicate the quantum level of that system. Krylov complexity is rooted in the Lanczos algorithm, a powerful method for assessing the nature of dynamical systems, which can be chaotic, integrable, or free. The key criteria for this assessment is the Lanczos coefficient, whose formula indicates the characteristics of a given dynamical system. For example, the relation $b=\alpha n$ suggests that the system is chaotic. In our paper, we highlight that the Lanczos coefficient effectively reveals that our early universe was a many-body, infinite, and chaotic system. From a cosmological perspective, the universe is considered an entirely open system. Notably, the Lanczos algorithm offers a natural way to construct the wave function of an open system using the second kind of Meixner polynomials. Given these insights, it is particularly important to explore Krylov complexity throughout the early universe, including inflation, RD and MD.
@@
In this paper, we thoroughly examine Krylov complexity in both open and closed system methods. Our primary result is the derivation of the two-mode squeezed state using the open system approach, along with the calculation of its evolution equations for $\phi_{k}(\eta)$ and $r_{k}(\eta)$. Below, we present our key findings.
$(a)$. In Sec. 2, we introduce the Lanczos algorithm and the associated Lanczos coefficients, which apply equally to both open and closed systems. As noted by Parker:2018yvk , the Liouvillian superoperator acts as the Hamiltonian when it can be expressed in terms of creation and annihilation operators. Our exploration of Krylov complexity encompasses the entire universe, including inflation, RD, MD, and the violations of slow-roll conditions, which necessitates considering contributions from potential energy. Sec. 3 discusses the most common potentials, such as the Higgs potential, Starobinsky’s inflationary potential, and chaotic inflationary potential. These models encompass a wide range of inflationary scenarios due to the unifying characteristics of $\alpha$ attractors. To accurately capture the information pertaining to these potentials, we examine the dynamics of the field $\phi$ in relation to the scale factor during the preheating process, as represented in Eq. (21). From these calculations, we determine the effective mass for various inflationary potentials, as illustrated in Fig. 2. Our findings indicate that the potential approaches certain constant values during RD and MD.
@@
$(b)$. In light of the scale factor and various potentials, we examine Krylov complexity using the closed system method. First, we numerically simulate the $r_{k}$ of the two-mode squeezed state, as depicted in Fig. 3. The simulation shows oscillations during inflation, which are enhanced in RD and MD, maintaining some constant values. Notably, the different potentials produce almost identical trends for $r_{k}$. Based on the numerical results for $r_{k}$, we also simulate the corresponding Krylov complexity and Krylov entropy, as illustrated in Figs. 4 and 5. The trends of both Krylov complexity and Krylov entropy are similar across the corresponding epochs. During inflation, we observe exponential growth, consistent with our previous study Li:2024kfm . In the RD and MD phases, the trends show initial enhancement before saturating to constant values. Additionally, various inflationary models exhibit nearly the same trends for Krylov complexity and Krylov entropy, confirming the approximation that the effective mass of the inflaton is approximately constant, as indicated in Li:2024ljz .
@@
$(c)$. Within the method of an open system, we rigorously derive the wave function of an open system using the second kind of Meixner polynomials, which we name it as the open two-mode squeezed state. This derivation is represented in Eq. (73). Throughout this process, we discover that the Lanczos coefficient $b_{n}$ and the dissipative coefficient are precisely determined by the specific Hamiltonian, leading to the conclusion that $b_{n}$ is identical across both methods. Our findings indicate that $b_{n}=\alpha n$, suggesting that our universe behaves as an entirely chaotic system. Consequently, the Lyapunov index is equivalent to $b_{n}$ when $n=2$. Figure 6 illustrates the chaos present across various epochs. During the inflationary period, chaos increases exponentially before diminishing to a certain level, only to rise again due to the preheating process. Additionally, we examine the evolution of the dissipative coefficient in RD and MD. This analysis builds on previous work in Ref. Li:2024kfm , which discusses the strong dissipative nature of the system during inflation. From Fig. 7, we can confirm that the RD and MD systems are weakly dissipative, characterized by $\mu_{2}\ll 1$.
Additionally, we derive the evolution of $r_{k}$ and $\phi_{k}$, noting that the solution for $r_{k}$ is dependent on the solution for $\phi_{k}$ due to their coupling, as described by Eq. (74). Using the wave function in Eq. (73), we calculate the Krylov complexity and Krylov entropy, which are represented in Eqs. (76) and (77). The leading order of these quantities aligns well with the formulas derived from the closed system method. In accordance with standard procedures, we need the numerical values of $r_{k}$ within the open system framework, as illustrated in Fig. 8. Unlike the closed system, there is no oscillation observed for $r_{k}$ during inflation, and the trends of Krylov complexity and Krylov entropy are depicted in Figs. 9 and 10. Compared to the closed system method, we observe that the values of Krylov complexity and Krylov entropy are significantly smaller. Recalling the definition of Krylov complexity, which pertains to the operator growth in a dynamical system, we find that a smaller operator growth suggests a decoherence-like behavior within the Krylov complexity. When dissipative effects are considered, we observe decoherence-like behavior, whether the system is weakly dissipative or not.
While we have achieved a comprehensive understanding of Krylov complexity in the early universe, many ideas still remain to be explored. Here, we outline several future directions.
Firstly, we investigate the Krylov complexity in the context of single-field inflation, and it is quite natural to extend our work into the multi-field framework Liu:2020zzv ; Liu:2019xhn ; Liu:2021rgq ; Zhang:2022bde as well as $f(R)$ gravity Liu:2018hno ; Liu:2020zlr ; Liu:2018htf . This extension allows us to explore the geometrical effects of fields and higher orders of Ricci curvature on Krylov complexity. Our Hamiltonian, given by (33), is Hermitian and includes the component of an open system. The traditional approach to handling open systems in quantum mechanics involves constructing what is known as a Lindbladian Lindblad:1975ef ; Gorini:1975nb , where the corresponding Hamiltonian is non-Hermitian. In our case, a natural method is to expand the action to higher orders with respect to the Mukhanov-Sasaki variable.
Secondly, our results indicate that the Krylov complexity in a weak dissipative system (RD or MD) will not follow the general discussions mentioned in Sec. 6.2. One possible explanation for this discrepancy is the neglect of decoherence in curvature perturbation during horizon exits. Ref. Bhattacharjee:2022lzy suggests that decoherence in circuit complexity may lead to this phenomenon of saturation. A similar concept might apply to Krylov complexity, particularly concerning the role of decoherence in the construction of the wave function.
@@
Thirdly, we have already obtained the open two-mode squeezed state as described in (73). In light this wave function, we can explore several additional avenues. For instance, we can calculate the two-point correlation functions to derive the power spectrum. Since the derivation of this wave function is model-independent, we have the flexibility to utilize different models of this wave function to test its validity. Conversely, we can also constrain the parameters $\mu_{1}$ and $\mu_{2}$ based on data from the Cosmic Microwave Background (CMB). Furthermore, it is important to highlight that Ref. Adhikari:2022oxr indicates the potential application of the CV conjecture to Krylov complexity during inflation; however, their approach is tailored for a closed system. In contrast, our universe is fundamentally an open system, which provides us with the opportunity to further investigate the validity of the CV conjecture for Krylov complexity within the context of open systems.
Acknowledgements.
@@
Appendix A The calculation of $r_{k}$ and $\phi_{k}$ within the method of open system
In this appendix, we will give the detailed calculations for $r_{k}$ and $\phi_{k}$ based on the open two-mode squeezed state,
@@ Appendix A The calculation of $r_{k}$ and $\phi_{k}$ within the method of open system
[…]
@@
[…]
@@
and therefore, we could obtain the evolution of $r_{k}$ as,
@@
the real component can read as,
@@
Finally, we could obtain the evolution equation of $\phi_{k}$ as follows,
@@
One can easily check that Eqs. (85) and (88) could nicely recover into Eq. (35) as $\mu_{1}=\mu_{2}=0$.

### v3

@@ VI.5 The evolution of $r_{k}(\eta)$ and $\phi_{k}(\eta)$ in open-system methodology
where prime denotes the varying with respect to the conformal time $\eta$. We change into the variable $y=\log_{10}a$, the evolution equation of $r_{k}$ and $\phi_{k}$ will become as,
@@
During the calculations, we accounted for the background of $\phi$ using conformal time $\eta$ via preheating, as demonstrated by the various potentials in Eq. (24). Consequently, $\mu_{2}$ and $\mu_{1}$ also vary with respect to conformal time, since they include contributions from the potential and the scale factor $a$.
@@
Fig. 8 clearly illustrates the evolution of $r_{k}$ across various epochs, including inflation, RD, and MD. It is evident that the different potentials produce similar trends for $r_{k}$. Unlike in Fig. 3, there is no oscillation of $r_{k}$ once dissipation effects are considered. Additionally, the trend of $r_{k}$ varies, particularly during the MD phase, where the evolution consistently shows an increase. These differences will result in distinctive evolutions of Krylov complexity and Krylov entropy.
VI.6 Krylov complexity and Krylov entropy with open-system methodology
In this section, we continue our investigation using the wave function (83) to calculate the Krylov complexity and Krylov entropy. The procedure for these calculations follows precisely that of our previous study Li:2024kfm . The only modification is the substitution of the parameter $r_{k}$ for $\eta$ in the formulas for Krylov complexity and Krylov entropy, which are given as follows:
@@ VI.6 Krylov complexity and Krylov entropy with open-system methodology
where we have defined $A=1+2\mu_{2}\tanh r_{k}+(\mu_{2}^{2}-|1-\mu_{1}^{2}|)\tanh^{2}r_{k}$. It straightforwardly demonstrates the leading order of Krylov complexity (86) and Krylov entropy (87) will reduce to Eq. (42) and Eq. (43) under the weak dissipative approximation, namely, $\mu_{2}\ll 1$.
Fig. 9 illustrates the evolution of Krylov complexity during inflation, RD, and MD. During inflation, Krylov complexity exhibits exponential growth, a behavior consistent with the closed-system methodology. In contrast, its evolution in the RD and MD epochs differs significantly with the closed-system methodology as shown in Fig. 4. For the open-system methodology, Krylov complexity in RD shows a decreasing trend, while in MD it first peaks as the scale factor $a$ increases and subsequently declines to a constant value. Notably, the magnitude of Krylov complexity in the open-system methodology is substantially lower than in the closed-system methodology across both RD and MD.
Our numerical analysis indicates that this suppression originates from the $\rm sech(r_{k})$ term in Eq. (86), an effect presents the dissipation strength. We have shown that both the RD and MD epochs can be approximated as weakly dissipative systems, characterized by $\mu_{2}<1$. Recalling that Krylov complexity quantifies operator growth in a many-body system, and our earlier finding $b_{n}=\alpha n$ indicates that the unverse is fully chaotic, the significantly reduced operator growth observed in the open-system methodology reveals that: dissipation generically leads to a faster suppression of operator growth compared to that in a closed-system methodology.
An understanding of the role of dissipation makes it clear that the magnitude of the Krylov entropy in open-system methodology is lower than that in the closed-system methodology, as demonstrated by a comparison of Figs. 10 and 5. The evolutionary trend is particularly distinct during the RD and MD epochs, as shown in Fig. 10. In RD, the Krylov entropy initially rises to a peak before saturating to a constant value. In MD, it first remains at a nearly constant level before declining and settling at a lower asymptotic value. Overall, both the cosmological background conditions and dissipative effects significantly influence the behavior of Krylov complexity and Krylov entropy.
@@
In this section, we employed the wave function (83) to study the Krylov complexity and Krylov entropy. By comparing our results with those from a closed-system methodology, we observed that dissipative effects not only alter the overall evolution of the system but also suppress the magnitude of both the Krylov complexity and entropy. Since Krylov complexity quantifies operator growth, a fundamental quantum feature of dynamical systems, the suppressed growth in the open-system methodology indicates that dissipation induces more rapid decoherence-like behavior. Furthermore, we note that different inflationary potentials yield nearly identical trends for both Krylov complexity and entropy across the various cosmological epochs.
@@ VII Summary and outlook
The definition of Krylov complexity is precise and unambiguous, particularly in contrast to the various definitions of circuit complexity. Its uniqueness remove the need to introduce a parametric manifold to conceptualize complexity. Furthermore, Krylov complexity characterizes the growth of operator supported within a dynamical system, and its magnitude indicates the effective quantum size of that system. Krylov complexity is fundamentally based on the Lanczos algorithm, a powerful method for diagnosing the nature of dynamical systems, whether chaotic, integrable, or free. The key diagnostic is the behavior of the Lanczos coefficient $b_{n}$, whose asymptotic growth with index $n$ serves as an indicator of the system’s dynamical phase. For instance, a linear growth $b_{n}\sim\alpha n$ is a hallmark of quantum chaos. In our work, we emphasize that the Lanczos coefficient explicitly reveals the early universe as a many-body, infinite, and chaotic system. From a cosmological standpoint, the universe is fundamentally an open system. Notably, the Lanczos algorithm provides a natural framework for constructing the wave function of an open system using the second kind of Meixner polynomials. Given these insights, it is particularly important to investigate Krylov complexity throughout the evolution of the early universe, encompassing the inflationary, RD, and MD eras.
@@
In this paper, we thoroughly examine Krylov complexity in both open- and closed-system methodologies. Our primary result is the derivation of the two-mode squeezed state using the open-system methodology, along with the calculation of its evolution equations for $\phi_{k}(\eta)$ and $r_{k}(\eta)$. Below, we present our key findings.
$(a)$. In Sec. II, we introduce the Lanczos algorithm and the associated Lanczos coefficients, a framework applicable to both open and closed systems. As highlighted in Parker:2018yvk , the Liouvillian superoperator effectively acts as a Hamiltonian when expressed in terms of creation and annihilation operators. Our investigation of Krylov complexity spans the entire evolution of the universe, including the inflationary epoch, RD and MD eras, where the slow-roll conditions are violated, the latter necessitating the inclusion of potential energy contributions.
Sec. III examines several representative inflationary potentials, such as the Higgs potential, the Starobinsky inflationary potential, and chaotic inflationary models. These models collectively cover a broad ranges of inflationary scenarios, unified through the characteristics of $\alpha$-attractors. To accurately extract the dynamical information encoded in these potentials, we analyze the evolution of the field $\phi$ relative to the scale factor during preheating, governed by Eq. (22). From this analysis, we compute the effective mass for the various inflationary potentials, as shown in Fig. 2, where our results indicate that how the effective mass varies with respect to scale factor.
@@
$(b)$. Considering the scale factor and various potentials, we investigate Krylov complexity using the closed-system methodology. First, we numerically simulate the parameter $r_{k}$ for the two-mode squeezed state, as shown in Fig. 3. The results reveal oscillations before horizon exits and then it exponetially grows during inflation, which become enhanced in the RD and MD eras, eventually settling into constant values. Notably, different potentials yield nearly identical trends for $r_{k}$. Based on the numerical results for $r_{k}$, we also simulate the corresponding Krylov complexity and Krylov entropy, as depicted in Figs. 4 and 5. Both Krylov complexity and entropy exhibit similar trends across the respective epochs. During inflation, we observe exponential growth, consistent with our previous study Li:2024kfm . In the RD and MD phases, the trends display an initial enhancement followed by saturation to constant values. Moreover, various inflationary models demonstrate nearly identical trends for both Krylov complexity and entropy, confirming the approximation that the effective mass of the inflaton is approximately constant, as indicated in Li:2024ljz .
@@
$(c)$. Within the open-system methodology, we rigorously derive the wave function of an open system using the second kind of Meixner polynomials, dubbed as the open two-mode squeezed state. This wave function is represented in Eq. (83). Throughout this process, we discover that the Lanczos coefficient $b_{n}$ and the dissipative coefficient $\mu_{2}$ are precisely determined by the specific Hamiltonian, leading to the conclusion that $b_{n}$ is identical across both methodologies. Our findings indicate that $b_{n}=\alpha n$, suggesting that our universe behaves as an entirely chaotic system. Consequently, the Lyapunov index is equivalent to $b_{n}$ when $n=2$. Figure 6 illustrates the chaos present across various epochs. During the inflationary period, chaos increases exponentially before diminishing to a certain level, only to rise again due to the preheating process. Additionally, we examine the evolution of the dissipative coefficient in RD and MD. This analysis builds on previous work in Ref. Li:2024kfm , which discusses the strong dissipative nature of the system during inflation. From Fig. 7, we confirm that the RD and MD systems are weakly dissipative, characterized by $\mu_{2}\ll 1$.
Additionally, we derive the evolution of the squeezing parameters $r_{k}$ and $\phi_{k}$. Their solutions are coupled, as described by Eq. (84), meaning the solution for $r_{k}$ depends on that for $\phi_{k}$. Using the wave function given in Eq. (83), we compute the Krylov complexity and Krylov entropy, presented in Eqs. (86) and (87). Their leading-order forms are in good agreement with the corresponding formulas derived within the closed-system methodology.
Following the standard procedure, we require the numerical values of $r_{k}$ in the open-system framework, as shown in Fig. 8. In contrast to the closed-system case, no oscillations in $r_{k}$ are observed during inflation. The resulting trends for Krylov complexity and Krylov entropy are depicted in Figs. 9 and 10, respectively. Compared to the closed-system results, the magnitudes of both Krylov complexity and entropy are significantly reduced. Recalling that Krylov complexity quantifies operator growth in a dynamical system, a smaller value indicates suppressed operator growth. This suppression reflects a decoherence-like behavior in the context of Krylov complexity once considering the open-system methodology. While we have achieved a comprehensive understanding of Krylov complexity in the early universe, many ideas still remain to be explored. Here, we outline several future directions.
Our investigation begins with Krylov complexity in single-field inflation. A natural extension of this work involves generalizing it to multi-field frameworks Liu:2020zzv ; Liu:2019xhn ; Liu:2021rgq ; Zhang:2022bde and to theories of $f(R)$ gravity Liu:2018hno ; Liu:2020zlr ; Liu:2018htf . Such extensions would allow us to explore the influence of field-space geometry and higher-order Ricci curvature terms on Krylov complexity. Our analysis is based on the Hamiltonian given in (35), which is Hermitian and incorporates open-system methodology. In conventional quantum mechanics, open systems are often treated using a Lindbladian master equation Lindblad:1975ef ; Gorini:1975nb , typically involving a non-Hermitian effective Hamiltonian. In our present framework, a systematic approach involves expanding the action to higher orders in the Mukhanov-Sasaki variable.
Secondly, our findings demonstrate that in a weakly dissipative system (such as during RD or MD), the behavior of Krylov complexity deviates from the general behavior outlined in Sec. VI.2. One plausible explanation for this discrepancy is the neglect of decoherence effects in the curvature perturbations upon horizon exit. As suggested in Ref. Bhattacharjee:2022lzy , decoherence can lead to a saturation phenomenon in circuit complexity. A conceptually similar mechanism may govern Krylov complexity, particularly regarding the influence of decoherence on the structure of the wave function.
@@
Thirdly, having derived the open two-mode squeezed state wave function presented in (83), several promising avenues for further investigation emerge. For instance, this wave function can be used to calculate two-point correlation functions and derive the corresponding power spectrum. Given that its derivation is model-independent, it offers the flexibility to test its validity by applying it within different theoretical models. Conversely, parameters such as $\mu_{1}$ and $\mu_{2}$ can be constrained using observational data from the Cosmic Microwave Background (CMB). Furthermore, it is noteworthy that Ref. Adhikari:2022oxr suggests the potential applicability of the Complexity=Volume (CV) conjecture to Krylov complexity during inflation, although their analysis is confined to a closed system. In contrast, since our universe is fundamentally an open system, our framework provides a distinct opportunity to investigate the validity of the CV conjecture for Krylov complexity within an open-system methodology.
Acknowledgements
@@
Appendix A The calculation of $r_{k}$ and $\phi_{k}$ within open-system methodology
In this appendix, we present the detailed derivations and calculations for the parameters $r_{k}$ and $\phi_{k}$ based on the open two-mode squeezed state,
@@ Appendix A The calculation of $r_{k}$ and $\phi_{k}$ within open-system methodology
Starting from the Schrödinger equation $\hat{H}\psi=i\partial_{\eta}\psi$ (with $\hbar=1$), we first evaluate its left-hand side as follows.
@@
Substituting Eqs. (90) and (91) into the Schrödinger equation yields the following simplified form for the open system’s squeezed state:
@@
and therefore, we obtain the evolution of $r_{k}$ as,
@@
the real component reads as,
@@
Finally, we obtain the evolution equation of $\phi_{k}$ as follows,
@@
It can be verified that Eqs. (95) and (98) correctly reduce to Eq. (37) in the limit $\mu_{1}=\mu_{2}=0$.